OpenStudy (shootforthestars):

Solve a system to find the ocean depth d1.

OpenStudy (shootforthestars):



OpenStudy (shootforthestars):

d1=.14*V d2=.223*V d2=sqrt((d1^2)+(62500))

OpenStudy (unklerhaukus):

d1 is given?

OpenStudy (unklerhaukus):

what is V ?

OpenStudy (shootforthestars):

That is in terms of velocity. I need to find the length

OpenStudy (unklerhaukus):

i am having trouble interpreting the digram, are there similar triangles or right angles?

OpenStudy (shootforthestars):

d2 is equal to d3

OpenStudy (unklerhaukus):